OpenStudy (anonymous):

Find the derivative of the function: y = e^(9 − x)

sam (.sam.):

you can start by taking the natural log of both sides, \[ y = e^{9 − x}\] \[\ln(y)=9-x\] Differentiate it \[\frac{1}{y}\frac{dy}{dx}=-1 \\ \\ \frac{dy}{dx}=-y\] Substitute y = e^(9 − x) into that will give you the derivative of y

OpenStudy (anonymous):

Thank you!!